I need help with this problem,
7-x/4=5(x+1)/9

Answer by rchill(405)

If you multiply both sides by the LCD, which is 36, you get 252-9x=20(x+1), which expands to 252-9x=20x+20. Now add 9x to both sides to get 252=29x+20, and then subtract 20 from both sides to get 232=29x. Finally divide both sides by 29 to get x=8.